OpenStudy (anonymous):

Find the value of each reciprocal function: Tan A=4.5, cot A=?

Nnesha (nnesha):

just flip the fraction example reciprocal of a/b is b/a

OpenStudy (anonymous):

So 5.4?

Nnesha (nnesha):

no...

Nnesha (nnesha):

\[\rm Tan \theta =\frac{ 1 }{ \cot\theta }\] 4.5 is same 4.5 over 1 now flip the fraction

Nnesha (nnesha):

\[\rm Tan A=4.5 =\frac{ 4.5 }{ 1 }\]

OpenStudy (anonymous):

1/4.5?

Nnesha (nnesha):

:=)

OpenStudy (anonymous):

So 1/4.5 is what Cot A equals?

Nnesha (nnesha):

yes

OpenStudy (anonymous):

Can you help me with a few more?

OpenStudy (anonymous):

Sin A=1/3, csc A=? Cot A=1.32, tan A=? Write the following in terms of the appropiate confunction: Csc 71 degrees Tan A Cot 23 degrees

Nnesha (nnesha):

there are reciprocal of each othr like \[\large\rm {sin \theta = \frac{ 1 }{ \csc \theta }~~~\cos \theta =\frac{ 1 }{ \sec \theta }~~\tan \theta =\frac{ 1 }{ \cot \theta} }\]

Nnesha (nnesha):

yea that's what you need ^^^

Nnesha (nnesha):

wait that's only one question right ?

OpenStudy (anonymous):

No it's 6 seperate ones.

Nnesha (nnesha):

hmm post one by one i don't get it. ..

OpenStudy (anonymous):

Sin A=1/3, csc=?

OpenStudy (anonymous):

Cot A=1.32, tan A=?

OpenStudy (anonymous):

Csc 71 degrees

OpenStudy (anonymous):

Tan A

Nnesha (nnesha):

alright wait

OpenStudy (anonymous):

Cot 23 degrees

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @briannababe Sin A=1/3, csc=? \(\color{blue}{\text{End of Quote}}\) what is reciprocal of sin ?

OpenStudy (anonymous):

Cosine?

Nnesha (nnesha):

sin = 1/csc csc = 1/ sin if sin = 1/3 then csc = what ( remember sin is reciprocal of csc )

OpenStudy (anonymous):

1/1/3

Nnesha (nnesha):

yes now multiply the numerator by the reciprocal of the bottom fraction

Nnesha (nnesha):

Or just flip the fraction if sin =a/b then csc = b/a

OpenStudy (anonymous):

3/1?

Nnesha (nnesha):

right!

Nnesha (nnesha):

3/1 is same as just 3

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @briannababe Cot A=1.32, tan A=? \(\color{blue}{\text{End of Quote}}\) same thing!

OpenStudy (anonymous):

For for the second problem which is cot A=1.32, tan A=? The answer would be 32.1? Or 32.1/1

Nnesha (nnesha):

wait no

Nnesha (nnesha):

cot A =1.32 what is reciprocal of 1.32 ?

Nnesha (nnesha):

it's 1.32 not 32.1

OpenStudy (anonymous):

Yeah lol i dont know how i typed 32.1 sorry

Nnesha (nnesha):

it's oky

OpenStudy (anonymous):

So the answer would be 1.32/1

Nnesha (nnesha):

now 1.32 = 1.32 but you need reciprocal of 1.32/1

Nnesha (nnesha):

here again \[\huge\rm if ~~\sin = a ~~~then~ \csc =\frac{ 1 }{ a }\]

OpenStudy (anonymous):

Ummm... I dont know

OpenStudy (anonymous):

Im lost

OpenStudy (anonymous):

Oh wait 1/1.32

Nnesha (nnesha):

you know ... just focus

Nnesha (nnesha):

yeah!!

OpenStudy (anonymous):

What about this it says write the following in terms of the appropiate confunction. Csc 71 degrees